OpenStudy (anonymous):

^3sqrt6y^8/^3sqrt125x^9 divide and multiply

jhonyy9 (jhonyy9):

y^2cuberoot6 ------------- 5x^3

jhonyy9 (jhonyy9):

^3sqrt is cuberoot

OpenStudy (anonymous):

\[\frac{\sqrt[3]{6y^8}}{\sqrt[3]{125x^9}}=\frac{y^2\sqrt[3]{6y^2}}{5x^3}\]